OpenStudy (anonymous):

Find the value of each logarithm using the change of base formula. Logbase5 625.

OpenStudy (anonymous):

\[\log _{5}625\]

OpenStudy (whpalmer4):

The change of base formula says that \[\log_a x= \frac{\log x}{\log a}\] \[\log_5 625 = \frac{\log 625}{\log 5} =\]

OpenStudy (whpalmer4):

You should get a nice round number and be able to check your work by raising 5 to that power.

OpenStudy (anonymous):

i got 4 , how would i check that ?

OpenStudy (whpalmer4):

\[5^4 = 5*5*5*5 = \]

OpenStudy (anonymous):

thanks !